OpenStudy (anonymous):

1. Management at a radio station wants to know the number of songs it plays per hour. A random sample over the past month yields the following counts for songs per hour: 9, 13, 12, 8, 7, 10, 11, 10, 8, 12, 13, 9, 10, 12, and 11.

OpenStudy (anonymous):

Ho: μ<10, Ha: μ>10

OpenStudy (anonymous):

test statistic (t test): t=(x̅-μ)/[s/√(n)]=(10.33-10)/(1.88/√(15))=.680

OpenStudy (anonymous):

what would be the p-value?

OpenStudy (anonymous):

i just need the steps.

OpenStudy (isaiah.feynman):